a box of mass

m = 12.6 kg

is released from rest (at position A) at the top of a 30.0
Answer #1

vb^2=2*gsin(theta)*d=2*9.8*0.5*4.4

vb=6.56 m/s

At point C

enitire KE of the system = spring potenial energy

0.5*k.0.28^2=12.6*9.81*(4.4+0.28)*sin(30)

k=7378.52 N/m